OpenStudy (anonymous):

How do I find the instantaneous velocity of 0.55s?

OpenStudy (stormfire1):

We'll need some more info :)

OpenStudy (anonymous):

Time (s) 0.0 0.20 0.30 0.40 0.50 0.60 Position (cm [W]) 0 25 75 75 75 0 The correct answer is 7.5 cm/s [E]

OpenStudy (anonymous):

but I tried: -> Vin = 37.5 cm -------- 5s = 7.5 cm/s {?]

OpenStudy (stormfire1):

Well, from 0.50 to 0.60s you went from 75cm to 0. 75cm/10 = 7.5cm/s

OpenStudy (stormfire1):

Of course, that's Vavg during that time period...which isn't exactly what they're asking but if you know the answer is 7.5, then maybe the question is worded wrong

OpenStudy (anonymous):

So what should I write?

OpenStudy (stormfire1):

75/10 = 7.5 cm/s

OpenStudy (anonymous):

you got the 75 from?

OpenStudy (anonymous):

0.50 s/

OpenStudy (stormfire1):

From your table. At 0.5s, the distance is 75cm

OpenStudy (anonymous):

ohhh, okay. And if I find the instantaneous velocity at time 0.40s, the answer is 0.0 m/s why is that?

OpenStudy (anonymous):

if you don't mind me asking.. D:

OpenStudy (stormfire1):

Because you can see from the table that the object is at the same position before and after 0.40s

OpenStudy (stormfire1):

So it is implied that it isn't moving at that point which = 0 cm/s

OpenStudy (anonymous):

ohh, okay.

OpenStudy (anonymous):

So when I find the time of 0.55s, it is basically 75 (because of 0.50s) - 0 (final velocity) / 10 s (because of .50 to 0.60s)?

OpenStudy (stormfire1):

Exactly

OpenStudy (anonymous):

Also.. if I want to find the time of 0.10s I subtract 25 cm - 0 / 10?

OpenStudy (stormfire1):

yes

OpenStudy (anonymous):

why so? :(

OpenStudy (stormfire1):

Ok, disregard what I just deleted :) Here's a better way to explain the 7.5 cm/s: At 0.50s your position is 75cm which is 0.75m At 0.60s your position is 0cm or 0 m 0.75/0.10 = 7.5cm/s

OpenStudy (stormfire1):

I shoulda remember to convert the distance to meters first...which is why it didn't look right at first. Anyway, the last message I posted is the correct way to demonstrate the answer.

OpenStudy (anonymous):

ohh..

OpenStudy (stormfire1):

Sorry for the confusion...

OpenStudy (anonymous):

It's okay. So first I convert it to cm = m?

OpenStudy (stormfire1):

Yes. I try to always convert things first to meters, seconds, kilograms, etc and when I don't, I get into trouble :)

OpenStudy (anonymous):

ohhh, okay. But why?

OpenStudy (stormfire1):

All of the kinematic equations, force formulas, etc...are based on the standard units of meters, seconds, kilograms, etc.

OpenStudy (anonymous):

ohh..

OpenStudy (stormfire1):

Here's some of the standard SI units which you should use (converting if you need to): Length - meter Mass - kilogram Time - second Electric current - ampere Thermodynamic temperature - kelvin Amount of substance - mole Luminous intensity - candela

OpenStudy (anonymous):

What if I want to find the instantaneous velocity at time 0.55 s, isn't that like the middle? of 0.50 - 0.60? Which will have the middle of 75 which is 37.5 cm?

OpenStudy (stormfire1):

one sec...this may help

OpenStudy (anonymous):

I'm so sorry for bothering you D;

OpenStudy (stormfire1):

It's no bother at all :) \[V_{inst}=\frac{\Delta d}{\Delta t}\] Where d is displacement and t is time. The deltas are always calculated as the final value minus the initial value. Anyway, what that means is that there's no need to worry about what's in the middle when calculating the delta. So the difference in displacement between 0.50s and 0.60s is 0m - 0.75m = -0.75m. The difference in time is 0.60s - 0.50s = 0.10s Note that the distance is negative because you're essentially heading back to 0 (to the left) on the x-axis

OpenStudy (stormfire1):

err..in the last line I meant "the velocity is negative"...I'm all screwed up today

OpenStudy (anonymous):

yup, I knew about the final and initial value. Really? I didn't know that I didn't have to know what's in the middle when calculating the delta.. But how come the correct answer is 7.5 cm [E]?

OpenStudy (stormfire1):

If you actually use the formula you get:\[V_{inst}=\frac{-0.75m}{0.1s}=-7.5 m/s\]It's magnitude is 7.5 and it's direction is left (since it's negative)

OpenStudy (anonymous):

yup. But I'm confused about the East since the data shows West.

OpenStudy (anonymous):

and the answer is 7.5 cm/s

OpenStudy (stormfire1):

The delta is merely the difference between the 2 values...what's in the middle doesn't really matter. In reality, you have no idea what's between those two values because you only know what the values are at those specific points. For all you know, the velocity have jumped up to 20 cm/s at 0.55s...but you have to go on what you do know.

OpenStudy (anonymous):

ohh, yeah.

OpenStudy (stormfire1):

Ok, I didn't see the West part. That explains why your final answer is positive...which I thought was just because they only wanted the magnitude. So the revised (and hopefully final equation) would be: \[V_{inst}=\frac{(0m - (-0.75m)}{(0.60s - 0.50s)} = +7.5 m/s\]

OpenStudy (anonymous):

why the -0.75 m?

OpenStudy (stormfire1):

I'm assuming West is negative:

OpenStudy (anonymous):

because it's going down?

OpenStudy (anonymous):

ohhh, yup.

OpenStudy (stormfire1):

I hope I havent confused you more by all this :) I have it straight in my head but it doesn't seem to be coming out very clear

OpenStudy (anonymous):

oh my, it just feels overwhelming because there are so much info but thank you so much for trying to explain these concepts to me! I really appreciate them

OpenStudy (stormfire1):

You'll get the hang of it...it just takes some practice. In reality, you can do this problem without converting units but once you move on to harder problems, it's safer just to convert them first and then work the problem. That's my opinion anyway

OpenStudy (anonymous):

yeah, I'll remember that! Thank you so much!

OpenStudy (anonymous):

btw, when I find the instantaneous velocity of the time: 0.10s which has 25 cm [W], I subtract

OpenStudy (stormfire1):

brb...wife

OpenStudy (anonymous):

it will be 0.25 m and 0-0.25 m / 10 s =

OpenStudy (anonymous):

ahh, okay.

OpenStudy (stormfire1):

Sorry for the delay...one sec